Write the equation in slope intercept form.
x+y-5=0

Answer by wgomero(28)

resolution:
Equation in slope intercept form: y = mx + b
equation x + y -5 = 0;
solving by y: y = -x + 5
need to find two points and the slope
Points:
if x=0 then y = 0 + 5; y=5; (0, 5)
if x=1 then y = -1 + 5; y=4; (-1,4)
slope m = (5 -4)/(0 -(-1)); 1/1 = 1; therefore m =1
we need to find b (y intercept)using point (0,5)
5 = 1(0) + b; b = 5
equation:
y = 1(x) + 5; y = x + 5